find the mean and the standard deviation of the heights of male engineering students in which the heights 66 inches and 75 inches correspond to standard scores of -0.6 and 1.1 respectively
1
Expert's answer
2020-11-09T19:52:58-0500

"Given \;that,\\
y_{1}=66,y_{2}=75,Z_{1}=-0.6,Z_{2}=1.1,then\\
Z=\frac{y-\mu}{\sigma}\\
\therefore -0.6=\frac{66-\mu}{\sigma}\implies -0.6\sigma=66-\mu,\\
\therefore 1.1=\frac{75-\mu}{\sigma}\implies 1.1\sigma=75-\mu,\\
\text{by multiplying the first equation by (-1) }\\
\text{and then adding the result to the second equation}\\
1.7 \sigma=-66+75=9\implies \sigma=\frac{9}{1.7}\approx 5.294,\\
\mu=66+0.6(5.294)=69.1764\\"